Hellen needs 20 000 in three years time.what is the least amount that she should deposit now into an account that offers an annual interest rate of 15%,compounded every 3 months,to have the required amount by then? The answer to the nearest rand is?
Answer by Theo(13342)

f = p * (1 + r) ^ n

f = future value
p = present value
r = interest rate per time period
n = number of time periods.

in your problem.

f = 20,000
r = 15% per year compounded quarterly (every 3 months)
n = 3 years.

use interest rate as a decimal rather than percent.
15% = .15
divide interest rate by number of compounding periods per year to get:
r = .15/4 = .0375 per quarter.

multiply number of years by 4 to get number of quarters.
3 * 4 = 12 quarters.

f = p * (1 + r) ^ n becomes:

20,000 = p * (1 + .0375) ^ 12

solve for p to get:

p = (20,000 / (1 + .0375) ^ 12).

this results in p = 12,857.97956

that's how much needs to be invested today in order to have 20,000 in 3 years at 15% per year compounded quarterly.

12,857.97956 * (1 + .0375) ^ 12 = 20,000